OpenStudy (anonymous):

Solve m= 2E/v^2.. solve for E

OpenStudy (anonymous):

Start by multiplying both sides by v^2, then what have you got?

OpenStudy (anonymous):

M=\[\frac{ 2E }{ v^2 } \] solve for e

OpenStudy (anonymous):

Yes, multiply both sides by v^2, first. Then what do you have?

OpenStudy (anonymous):

\[\frac{ M }{ v^2} = \frac{ 2e }{ v^2 }\]

OpenStudy (anonymous):

No, multiply by v^2. In order to dispel of a division, you must multiply. What you do to one side of an equation you must also do to the other.

OpenStudy (anonymous):

Draw it out for me lol

OpenStudy (anonymous):